Consider the reaction

A—> B



The rate of the reaction is 1.6 x 10 ^-2 m/s when the concentration of A is 0.35 M. Calculate the rate constant if the reaction is (a) first order in A and (b) second order in A.



1
Expert's answer
2021-10-15T12:33:19-0400

(a) First order reaction:

Rate = k(A),


Where (k) is a constant


R = [ ( 1.6)(10-2)](0.35)


R = (5.6)(10-2)ms-1


R = 0.056ms-1


(b) Second order reaction:


Rate = k ( A) 2


R = [(1.6)(10)-2 ](0.35)2


R = (1.96)(10-3)m-1s-1


R = 0.00196m-1s-1-